# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v1

@@ 3.3 Data Analysis
Fig. 5(a) shows the distribution of objects across five categories in day and night conditions, with cars being the most common in V2X-Radar, followed by cyclists and pedestrians, while trucks and buses are the least frequent. Fig. 5(b) presents the maximum and average number of LiDAR points within 3D bounding boxes for each category, indicating that larger vehicles have more 3D points than smaller pedestrians or cyclists. Fig. 5(c) reveals the density distribution of 4D Radar points within different object’s bounding boxes, mirroring a similar trend to Fig. 5(b). Lastly, Fig. 5(d) demonstrates that annotations per collaborative sample can reach up to 90 - significantly higher than single-vehicle datasets like KITTI [5] or nuScenes [1], highlighting how combining vehicle-side and roadside data enhances environmental perception comprehensiveness.
@@ 4 Tasks
Our V2X-Radar dataset, comprising of three subsets: V2X-Radar-I, V2X-Radar-V, and V2X-Radar-C, is designed to facilitate roadside 3D object detection, single-vehicle 3D object detection, and cooperative 3D object detection.
@@ 4.1 Single-agent 3D Object Detection
Task Definition. Single-agent 3D object detection involves two tasks: roadside 3D object detection using the V2X-Radar-I sub-dataset and vehicle-side 3D object detection using the V2X-Radar-V sub-dataset. These tasks rely on sensors from either the intelligent roadside unit or the vehicle-side platform for detecting 3D objects, leading to the following challenges:
  • Single-modal Encoding: The process of encoding 2D image from the camera, dense 3D point cloud from LiDAR, sparse point cloud with Doppler information from 4D Radar into 3D spatial representation is necessary for precise single-modal 3D object detection.
  • Multi-modal Fusion: When fusing multi-modal information from different sensors, it is necessary to consider (1) spatial misalignment, (2) temporal misalignment, and (3) sensor failure. These issues are necessary to achieve robust multi-modal 3D object detection.
Evaluation Metrics. The evaluation area covers [-100, 100] in the x direction and [0, 100] in the y direction from the ego-vehicle or roadside unit. Following the metrics in Rope3D [40], KITTI [5] datasets, we utilize the Average Precision (AP) at IoU thresholds of 0.5 and 0.7 to evaluate the performance of object detection.
Benchmark Methods. We extensively evaluate various leading single-agent 3D object detectors based on different sensor inputs. In particular, our evaluation encompasses LiDAR-centric techniques such as PointPillars [9], SECOND [36], CenterPoint [41], and PV-RCNN [19]; camera-dependent methods including SMOKE [13], BVDepth [11], BEVHeight [38], and BEVHeight++ [37]; and 4D Radar-based approaches like RPFA-Net [29] and RDIoU [18].
Table 3: Roadside 3D object detection benchmarks on V2X-Radar-I under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, R denotes LiDAR, Camera and 4D Radar respectively.
@@
Table 4: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, R denotes LiDAR, Camera and 4D Radar respectively.
@@ 4.2 Cooperative 3D Object Detection
Task Definition. The goal of the cooperative 3D object detection task using the V2X-Radar-C sub-dataset is to utilize sensors from both the vehicle and intelligent roadside unit for ego-vehicle’s 3D object detection. Unlike previous single-agent perception, cooperative perception presents unique domain-specific challenges:
  • Spatial Asynchrony: Localization errors can create discrepancies in the relative position between the single-vehicle platform and the intelligent roadside unit, potentially leading to global misalignment when translating data from the roadside unit to the ego-vehicle coordinate system.
  • Temporal Asynchrony: Communication delays in data transmission can result in timestamp discrepancies between sensor data from the single-vehicle platform and the intelligent roadside unit. This may cause local misalignment of dynamic objects when translating data within the same coordinate system.
Evaluation Metrics. The evaluation area covers [-100, 100] meters in both the x and y directions with respect to the ego vehicle. Similar to DAIR-V2X [42] and V2V4Real [35], We group various vehicle types into the same class and focus solely on vehicle detection. The object detection performance is evaluated by Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are measured through Average MegaBytes (AM). Following prior studies [42, 35], we compare methods in two setups: (1) Synchronous, ignoring communication delays. (2) Asynchronous, with a 100 ms delay simulated by fetching the roadside sensor data with the previous timestamp.
Benchmark Methods. We present comprehensive benchmarks for the following three fusion strategies in cooperative 3D object detection:
  • Late Fusion: Each agent utilizes its sensors to detect 3D objects and shares the predictions. The receiving agent then applies NMS to produce the final outputs.
  • Early Fusion: The ego-vehicle gathers all the point clouds both from itself and other agents to its own coordinate system, then proceeds with detection procedures.
  • Intermediate Fusion: Each agent uses a neural feature extractor to obtain intermediate features, which are then compressed and sent to the ego vehicle for cooperative feature fusion. We evaluate several leading methods, including F-Cooper [2], V2XVIT [32], CoAlign [31], and HEAL [16], to establish a benchmark in this field.
@@ 5.1 Implementation details
For the dataset split, single-agent 3D object detection, including roadside and vehicle-side scenarios, adheres to two distinct setups: (1) homogeneous split, the dataset is divided into train/val/test with 7000, 1500, and 1500 frames, respectively. It is crucial to highlight that the selection of samples in this setup is completely random. (2) heterogeneous split, the dataset is divided into train/val/test with 34, 8, and 8 sequences, respectively. The dataset for cooperative 3D object detection is divided into train/val/test with 30, 5, and 5 sequences, respectively. In particular, all experimental results are evaluated using the validation set. For the detection areas, the single-agent detection techniques span a range of [0, 100]m in the x-direction and [-100, 100]m in the y-direction relative to the ego vehicle or roadside unit. On the other hand, cooperative 3D object detection encompasses a broader area, covering [-100, 100]m in both the x and y directions. For the ground truth, LiDAR or camera methods use all the annotations. Considering the limited coverage area of 4D Radar, 4D Radar-based methods only use labels located within its field of view. For the feature extractor, all the cooperative detection methods rely on PointPillar [9] to extract BEV features from LiDAR or 4D Radar point clouds and utilize BEVDet [7] to extract BEV features from camera images.
@@ 5.2 Benchmark results
Single-agent 3D Object Detection. The benchmark results for the V2X-Radar-I and V2X-Radar-V sub-datasets under a homogeneous split are presented in Tab. 3 and Tab. 4. Additional results under the heterogeneous split can be found in the supplementary materials. The experimental results in these tables clearly illustrate that LiDAR-based methods achieve the most superior performance. 4D radar-based methods, although they deal with a relatively sparse point cloud, still perform better than camera-based methods. Camera-based methods, hindered by their inability to utilize depth information, are less effective than LiDAR and 4D Radar-based methods.
Cooperative 3D Object Detection. Tab. 5 demonstrates a quantitative comparison among several methods on V2X-Radar-C. We can observe that:
  • Compared to the baseline of single-vehicle perception, all methods involving cooperative perception exhibit a significant performance improvement, highlighting its key role in boosting single-vehicle perception.
  • Compared to the synchronous setting results, introducing communication delay in asynchronous settings caused a significant drop in Average Precision (AP) across all methods. For example, with a strict IoU of 0.7, F-Cooper [2], CoAlign [15], and HEAL [16] using LiDAR point clouds experienced substantial decreases of 13.30%, 20.49%, and 18.51%, respectively. These findings underscore the necessity of mitigating the impact of communication delay in robust cooperative perception.
@@ 6 Conclusion
In this paper, we introduce V2X-Radar, the first large real-world multi-modal dataset featuring 4D Radar for cooperative perception, aiming to prompt cooperative perception research with 4D Radar for a broader operational design domain (ODD). Our dataset focuses on challenging intersection scenarios and provides data collected from diverse times and weather conditions. It comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data points, with 350K annotated bounding boxes spanning 5 categories. To facilitate a wide range of perception research, we structure the dataset into three sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we have conducted extensive benchmarks of recent perception algorithms on these three sub-datasets.
@@
Limitation and Future Work. Our dataset primarily focuses on 3D object detection, while the asynchronous setting in cooperative perception evaluation is simulated with a fixed time delay. In the future, we plan to expand our tasks based on V2X-Radar to include tracking and trajectory prediction, and we will explore performance evaluation under real communication delays of C-V2X devices.

### v2

@@ 3.3 Data Analysis
Fig. 5(a) illustrates the distribution of objects across five categories under both day and night conditions, with cars being the most prevalent in V2X-Radar, followed by cyclists and pedestrians, while trucks and buses are the least common. Fig. 5(b) displays the maximum and average number of LiDAR points within 3D bounding boxes for each category, suggesting that larger vehicles possess more 3D points than smaller pedestrians or cyclists. Fig. 5(c) reveals the density distribution of 4D Radar points within various objects’ bounding boxes, reflecting a trend similar to that shown in Fig. 5(b). Lastly, Fig. 5(d) indicates that annotations per collaborative sample can reach up to 90 - a significant increase compared to single-vehicle datasets such as KITTI [6] or nuScenes [1], emphasising how the integration of vehicle-side and roadside data enhances the comprehensiveness of environmental perception.
@@ 4 Tasks
V2X-Radar dataset comprises three sub-datasets: V2X-Radar-I, V2X-Radar-V, and V2X-Radar-C, which are designed for roadside 3D object detection, single-vehicle 3D object detection, and cooperative 3D object detection.
@@ 4.1 Single-agent 3D Object Detection
Task Definition. Single-agent 3D object detection involves two distinct tasks: roadside 3D object detection using the V2X-Radar-I sub-dataset, and vehicle-side 3D object detection employing the V2X-Radar-V sub-dataset. These tasks utilise sensors from either the intelligent roadside unit or the vehicle-side platform for 3D object detection, presenting the following challenges:
  • Single-modal Encoding: The process of encoding a 2D image from the camera, a dense 3D point cloud from LiDAR, and a sparse point cloud with Doppler information from 4D Radar into a 3D spatial representation is essential for precise single-modal 3D object detection.
  • Multi-modal Fusion: When fusing multi-modal information from various sensors, it is essential to account for (1) spatial misalignment, (2) temporal misalignment, and (3) sensor failure. Addressing these issues is crucial for achieving robust multi-modal 3D object detection.
Evaluation Metrics. The evaluation area covers [-100, 100] in x direction and [0, 100] in y direction from the ego-vehicle or roadside unit. Following the metrics in Rope3D [46], KITTI [6] datasets, we utilize the Average Precision (AP) at IoU thresholds of 0.5 and 0.7 to evaluate the performance of object detection.
Benchmark Methods. We conduct a comprehensive evaluation of various leading single-agent 3D object detectors based on different sensor inputs. Specifically, our evaluation encompasses LiDAR-centric techniques such as PointPillars [11], SECOND [40], CenterPoint [47], and PV-RCNN [21]; camera-dependent methods including SMOKE [15], BVDepth [13], BEVHeight [42], and BEVHeight++ [41]; and 4D Radar-based approaches like RPFA-Net [33] and RDIoU [20].
Table 3: Roadside 3D object detection benchmarks on V2X-Radar-I under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
@@
Table 4: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under homogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
@@ 4.2 Cooperative 3D Object Detection
Task Definition. The cooperative 3D object detection task with the V2X-Radar-C sub-dataset aims to utilise sensors from both the vehicle-side platform and the intelligent roadside unit to execute 3D object detection for the ego-vehicle. In contrast to the previous single-agent perception, cooperative perception presents domain-specific challenges:
  • Spatial Asynchrony: Localization errors can create discrepancies in the relative pose between the single-vehicle and the intelligent roadside unit, potentially leading to global misalignment when translating data from the roadside unit to the ego-vehicle coordinate systems.
  • Temporal Asynchrony: Communication delays in the data transmission process can result in timestamp discrepancies between the sensor data from the single-vehicle platform and the intelligent roadside unit. This may cause local misalignment of dynamic objects when translating data within the same coordinate systems.
Evaluation Metrics. The evaluation area spans [-100, 100] metres in both the x and y directions relative to the ego vehicle. Similar to DAIR-V2X [48] and V2V4Real [39], we group various vehicle types into the same class and concentrate solely on vehicle detection. The performance of object detection is evaluated using Average Precision (AP) at IoU thresholds of 0.5 and 0.7. Transmission costs are calculated through Average MegaBytes (AM). In line with previous studies [48, 39], we compare methods in two configurations: (1) Synchronous, ignoring communication delays. (2) Asynchronous, simulating data transmission delay by retrieving roadside samples with the preceding timestamp.
Benchmark Methods. We provide comprehensive benchmarks for the following three fusion strategies in cooperative 3D object detection:
  • Late Fusion: Each agent employs its sensors to detect 3D objects and shares the predictions. The receiving agent then applies NMS to generate the final outputs.
  • Early Fusion: The ego vehicle collects all point clouds from itself and other agents into its own coordinate system, then proceeds with detection procedures.
  • Intermediate Fusion: Each agent employs a neural feature extractor to obtain intermediate features; the encoded features are compressed and transmitted to the ego vehicle for cooperative feature fusion. We evaluate several prominent intermediate methods, including F-Cooper [2], V2XVIT [36], CoAlign [35], and HEAL [18], to establish a benchmark in this field.
@@ 5.1 Implementation details
For the dataset split, single-agent 3D object detection, which includes roadside and vehicle-side scenarios, follows two distinct setups: (1) homogeneous split, where the dataset is divided into train/val/test sets containing 7000, 1500, and 1500 frames, respectively. The selection of samples in this setup is entirely random. (2) heterogeneous split, where the dataset is divided into train/val/test sets with 34, 8, and 8 sequences. The cooperative 3D object detection dataset is divided into train/val/test sets with 30, 5, and 5 sequences. Notably, all experimental results are assessed using the validation set. For the detection areas, the single-agent detection techniques cover a range of [0, 100] m in the x-direction and [-100, 100] m in the y-direction relative to the ego vehicle or roadside unit. In contrast, cooperative 3D object detection encompasses a broader area, extending from [-100, 100] m in both the x and y directions relative to the ego vehicle. For the ground truth, LiDAR or camera methods utilize all the annotations. Considering the limited coverage area of 4D Radar, 4D Radar-based methods only use labels within its field of view for training and evaluation. For the feature extractor, all cooperative detection methods depend on PointPillar [11] to extract BEV features from LiDAR or 4D Radar point clouds and use LSS [9] to extract BEV features from camera images.
Figure 6: Cooperative 3D object detection benchmarks on V2X-Radar-C under different transmission delay.
@@ 5.2 Benchmark results
Single-agent 3D Object Detection. The benchmark results for the V2X-Radar-I and V2X-Radar-V sub-datasets under a homogeneous split are detailed in Tab. 3 and Tab.  4. Further results under the heterogeneous split can be found in the supplementary materials. The experimental results in these tables clearly demonstrate that LiDAR-based methods achieve the highest performance. Although 4D radar-based methods manage a relatively sparse point cloud, they still outperform camera-based methods. Camera-based methods, restricted by their inability to utilize depth information, are less effective than LiDAR and 4D Radar-based methods.
Cooperative 3D Object Detection. Tab. 5 presents a quantitative comparison of several methods on V2X-Radar-C:
  • In comparison to the benchmark of single-vehicle perception, all methods involving cooperative perception exhibit a significant performance enhancement, highlighting its essential role in improving single-vehicle perception. Compared to the results from the Sync setting, the introduction of communication delay in Async settings led to a considerable performance decline. As shown in Tab. 5 and Fig. 6, when subjected to a 100ms transmission delay with a strict 0.7 IoU threshold, F-Cooper [2], CoAlign [17], and HEAL [18] using LiDAR point cloud experienced significant reductions of 13.30%, 20.49%, and 18.51% respectively. These findings underline the necessity of mitigating the effects of communication delay to ensure effective and robust cooperative perception.
@@ 6 Conclusion
In this paper, we introduce V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar for cooperative perception, aiming to prompt cooperative perception with 4D Radar for a broader operational design domain (ODD). Our dataset focuses on challenging intersection scenarios and provides data collected under various times and weather conditions. It consists of 20K LiDAR frames, 40K camera images, and 20K 4D Radar data points, along with 350K annotated bounding boxes 5 categories. To support a wide range of perception research, we organise the dataset into three sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we have conducted extensive benchmarks of recent perception algorithms on these three sub-datasets.
@@
Limitation and Future Work. Our dataset focuses on 3D object detection, and we simulate asynchronous cooperative perception with a fixed time delay. In the future, we plan to expand our V2X-Radar tasks to include tracking and trajectory prediction, and evaluate performance under real C-V2X communication delays.
Supplementary Material
Figure 7: The time synchronization solution for the connected vehicle-side platform and the intelligent roadside unit. Both the vehicle platform and roadside unit are equipped with a computer, a time synchronization box, a GPS/IMU system, and various sensors, including LiDAR, 4D radar, and cameras. The GPS/IMU systems in both the vehicle platform and roadside unit receive the same GNSS signals from GPS satellites.
Table 6: Roadside 3D object detection benchmarks on V2X-Radar-I under the heterogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
Pointpillars [11] L 66.19 / 86.11 61.28 / 78.82 61.28 / 78.82 33.32 / 49.80 27.79 / 41.28 27.79 / 41.28 64.87 / 74.11 56.55 / 64.95 56.55 / 64.95
SECOND [40] L 68.93 / 86.82 62.11 / 79.48 62.11 / 79.48 39.76 / 56.78 32.83 / 47.03 32.83 / 47.03 69.73 / 77.73 60.79 / 68.29 60.79 / 68.29
CenterPoint [47] L 70.87 / 87.62 65.62 / 80.22 65.63 / 80.22 57.66 / 71.58 47.29 / 59.34 47.29 / 59.34 73.08 / 78.15 65.46 / 70.27 65.46 / 70.27
PV-RCNN [21] L 82.08 / 92.27 74.75 / 84.81 74.75 / 84.81 67.33 / 80.39 56.72 / 66.69 56.72 / 66.69 76.59 / 82.04 67.11 / 72.51 67.11 / 72.51
SMOKE [15] C 3.04 / 16.87 2.99 / 16.64 2.99 / 16.63 0.05 / 0.09 0.04 / 0.07 0.04 / 0.07 1.12 / 8.25 1.07 / 7.47 1.07 / 7.39
BEVDepth [13] C 12.24 / 23.06 11.45 / 21.76 10.72 / 21.58 0.08 / 2.11 0.06 / 0.86 0.06 / 0.86 1.75 / 12.92 1.56 / 10.24 1.55 / 10.13
BEVHeight [42] C 12.94 / 24.41 11.61 / 23.55 11.40 / 23.35 1.01 / 3.11 0.08 / 1.29 0.08 / 1.24 3.31 / 18.14 2.01 / 12.24 2.01 / 12.22
BEVHeight++[41] C 14.07 / 25.12 13.57 / 24.81 13.25 / 24.62 1.39 / 4.87 0.13 / 1.63 0.13 / 1.62 4.20 / 19.69 2.57 / 12.81 2.57 / 12.78
RDIoU [20] R 4.31 / 25.74 3.85 / 24.24 3.85 / 24.24 1.55 / 8.64 1.16 / 6.96 1.16 / 6.96 1.94 / 10.91 1.55 / 9.65 1.55 / 9.65
RPFA-Net [33] R 4.49 / 28.39 4.18 / 26.39 4.18 / 26.39 1.76 / 7.82 1.32 / 7.04 1.32 / 7.04 2.36 / 12.61 1.95 / 10.94 1.95 / 10.94
Table 7: Single-vehicle 3D object detection benchmarks on V2X-Radar-V under the heterogeneous split. The vehicle category includes car, bus and truck. ‘M‘ means modality, and L, C, and R denote LiDAR, Camera, and 4D Radar, respectively.
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
Pointpillars [11] L 72.03 / 80.43 64.65 / 72.87 64.65 / 71.87 39.01/44.17 33.37 / 40.17 33.37 / 40.17 72.46 / 75.99 61.97 / 65.82 61.97 / 65.82
SECOND [40] L 75.13 / 83.37 66.06 / 76.08 66.06 / 76.08 41.63 / 46.39 35.34 / 42.69 35.34 / 42.69 75.17 / 80.84 63.07 / 69.71 63.07 / 69.71
CenterPoint [47] L 77.67 / 85.22 67.24 / 78.26 67.24 / 78.26 53.77 / 57.67 48.08 / 54.68 48.08 / 54.68 83.92 / 89.51 72.69 / 75.69 72.69 / 75.69
PV-RCNN [21] L 83.34 / 87.91 77.37 / 81.24 77.37 / 81.24 62.33 / 63.18 56.87 / 59.87 56.87 / 59.87 86.68 / 91.54 76.73 / 78.05 76.73 / 78.05
SMOKE [15] C 0.31 / 7.28 0.16 / 3.81 0.16 / 3.74 0.03 / 0.18 0.02 / 0.11 0.02 / 0.11 0.08 / 1.89 0.07 / 1.25 0.07 / 1.25
BEVDepth [13] C 1.23 / 9.67 0.58 / 5 .57 0.58 / 5.41 0.06 / 1.02 0.04 / 0.71 0.04 / 0.71 1.01 / 3.74 0.05 / 1.91 0.05 / 1.91
BEVHeight [42] C 1.12 / 9.79 0.56 / 4.89 0.53 / 4.43 0.05 / 0.93 0.04 / 0.69 0.04 / 0.69 0.09 / 3.34 0.05 / 1.89 0.05 / 1.89
BEVHeight++[41] C 1.79 / 14.26 0.62 / 7.61 0.58 / 6.87 0.07 / 1.83 0.06 / 1.16 0.06 / 1.16 0.13 / 3.91 0.07 / 2.31 0.07 / 2.28
RDIoU [20] R 3.77 / 34.43 3.48 / 19.03 3.37 / 17.74 0.04 / 0.75 0.02 / 0.62 0.02 / 0.62 0.88 / 8.94 0.29 / 4.15 0.29 / 4.15
RPFA-Net [33] R 4.23 / 35.49 3.81 / 19.44 3.62 / 18.63 0.04 / 0.95 0.03 / 0.81 0.03 / 0.81 1.24 / 11.01 0.57 / 5.49 0.57 / 5.49
Figure 8: Diverse corner cases for single-vehicle autonomous driving. Each case is depicted by roadside multi-view images on the left and the vehicle-side image on the right. We use a red circle to mark the occluded objects.
Figure 9: Data collection across different periods and various weather conditions. a) Different times of day, including daytime, dusk, and nighttime. b) Various weather conditions like sunshine, rain, fog, and snow.
A Appendix
This supplementary document is organized as follows:
  • We offer further details on the time synchronization of sensors on vehicle and roadside platforms in Appendix B.
  • We provide supplementary benchmark results, which include roadside 3D object detection on V2X-Radar-I sub-dataset and vehicle-side 3D object detection on V2X-Radar-V sub-dataset, both evaluated under heterogeneous splits in Appendix C.
  • We provide visualizations of different data collection scenarios, including corner cases for single-vehicle autonomous driving and a variety of time-of-day and weather conditions in Appendix D.
B Synchronization Details
Time synchronization enables the simultaneous sampling of data from the same scene by various sensors on both the vehicle platform and roadside units. For a collaborative perception dataset, achieving precise time alignment is critical—not only among the diverse sensors within a single platform (be it the vehicle or the roadside unit) but also across sensors spanning both platforms.
The solution for time synchronization in our V2X-Radar is illustrated in Fig. 7. Typically, the vehicle platform and roadside unit align their sensors using a synchronization box that leverages the Precision Time Protocol (PTP). This synchronization box integrates Pulse Per Second (PPS) signals and Time of Day (ToD) data sourced from the GPS/IMU system. The process entails generating PTP timing signals to synchronize industrial computers, which subsequently facilitate hardware-triggered synchronization for LiDAR and 4D Radar. The PPS signals act as hardware triggers for camera data acquisition, utilizing rising edge pulses to ensure alignment between cameras, LiDAR, and 4D Radar within the same platform. To achieve time synchronization between the vehicle platform and the roadside unit, both systems receive identical GNSS signals from GPS satellites, ensuring precise clock alignment of the GPS/IMU systems across the two platforms.
C More Benchmark Results
We present benchmark results for single-agent 3D object detection, which include roadside 3D object detection on the V2X-Radar-I sub-dataset and single-vehicle 3D object detection on the V2X-Radar-V sub-dataset. All methods are evaluated using the heterogeneous split.
C.1 Roadside 3D Object Detection
Tab. 6 presents a quantitative evaluation of multiple roadside 3D object detection models, leveraging our V2X-Radar-I sub-dataset under a heterogeneous split configuration. The results reveal that:
  • LiDAR-based methods achieve the best performance, while 4D Radar and camera-based methods are less effective than LiDAR. Furthermore, the performance of 4D Radar and camera-based methods is comparable; 4D Radar excels at IoU=0.5, while camera-based methods perform better at the stricter IoU=0.7.
  • Compared to the results in Tab.4, both camera-based and 4D Radar-based roadside methods under the heterogeneous split show a significant decline in performance, highlighting that scenario generalization presents a challenge for both methods.
C.2 Single-vehicle 3D Object Detection
Tab. 7 presents comparative experiments of various single-vehicle 3D object detection models utilizing our V2X-Radar-V sub-dataset under a heterogeneous split. It is evident from the table that:
  • LiDAR-based methods demonstrate the highest performance. 4D Radar-based methods outperform camera-based methods but do not match the effectiveness of LiDAR-based approaches, exhibiting slight deviations from the results in Tab. 6.
  • In comparison to the findings presented in Tab. 5, both camera-based and 4D Radar single-vehicle approaches under the heterogeneous split exhibit markedly reduced performance. This underscores the substantial difficulties these methods face in achieving robust generalization across diverse scenarios.
D Data Collection Scenarios
Our data collection encompasses various weather conditions and times of day, focusing on challenging corner cases for single-vehicle autonomous driving. Fig. 8 presents a visual depiction of various challenging corner cases, such as occluded vehicles at a public intersection, partially hidden cyclists at a public T-junction, concealed pedestrians within a restricted park environment, and obscured cyclists at a campus intersection. Fig. 9 illustrates the scenarios employed for data acquisition, spanning diverse times of day and meteorological conditions. As highlighted in Fig. 9(a), the temporal segments include morning daytime, afternoon daytime, dusk, and nighttime. As shown in Figure 9(b), the weather conditions range from clear sunny skies to fog, rain, and snow.